# Visibility-Aware Language Aggregation for Open-Vocabulary Segmentation in 3D Gaussian Splatting

### v1

@@ Abstract
Recently, distilling open-vocabulary language features from 2D images into 3D Gaussians has attracted significant attention. Although existing methods achieve impressive language-based interactions of 3D scenes, we observe two fundamental issues: background Gaussians contributing negligibly to a rendered pixel get the same feature as the dominant foreground ones, and multi-view inconsistencies due to view-specific noise in language embeddings. We introduce Visibility-Aware Language Aggregation (VALA), a lightweight yet effective method that computes marginal contributions for each ray and applies a visibility-aware gate to retain only visible Gaussians. Moreover, we propose a streaming weighted geometric median in cosine space to merge noisy multi-view features. Our method yields a robust, view-consistent language feature embedding in a fast and memory-efficient manner. VALA improves open-vocabulary localization and segmentation across reference datasets, consistently surpassing existing works.
@@ 1 Introduction
Understanding 3D scenes is essential for interacting with the environment in robotic navigation [2, 22], autonomous driving [7, 31], and augmented reality [16, 9]. Traditional approaches, however, are constrained to a fixed set of object categories defined at training time [26, 4, 34], limiting their applicability to open-world scenarios. Thanks to recent advances in vision-language models [29, 10], open-vocabulary methods [8, 41, 23] enable querying and interacting with the 3D scene through natural language, and recognizing unseen object categories without retraining.
While classical 3D understanding works operate on point clouds or meshes derived from 3D sensors, recent neural scene representations such as NeRFs [21] and 3D Gaussian Splatting (3DGS) [13] have emerged as a compelling alternative. They not only enable high-quality rendering from novel viewpoints but also facilitate semantic reasoning, as appearance and geometry are encoded jointly. Thus, open-vocabulary reasoning has recently been extended to rendered 3D scenes [14, 27], enabling new semantic interactions in 3D environments. Initially explored with NeRFs [14, 6], the efficiency and explicit nature of 3DGS simplified the integration of semantic features, determining its widespread adoption [27, 37, 12, 3].
@@
Despite numerous recent works [12, 11, 17, 33], we found that current methods suffer from a fundamental limitation: they propagate the 2D features to all Gaussians along a camera ray, regardless of the scene geometry and occlusions. So, features from foreground objects (e.g., a vase) are incorrectly assigned to background elements (e.g., the table or floor behind it), severely degrading open-vocabulary performance.
Furthermore, when projected to 3D, the 2D features are inconsistent across views. The same object’s feature may drift from different locations or viewpoints, a phenomenon referred to as semantic drift[14]. Existing approaches attempt cross-view consistency with 3D-consistent clusters and contrastive signals guided by SAM masks [37, 19, 17, 25]. However, they typically require long per-scene optimizations, and their reliance on noisy, view-dependent 2D cues often threatens the clusters.
@@
  • We propose a robust aggregation strategy of the 2D features using the streaming cosine median and thus improve the multi-view consistency.
@@ 2 Related works
Other methods first group 3D Gaussians or points into semantically meaningful clusters, typically corresponding to objects or parts, and then assign a language feature to each cluster as a whole [19, 37, 28, 17, 25, 11]. These methods introduce an explicit discrete grouping step as a form of prior semantic structuring: OpenGaussian [37] performs coarse-to-fine clustering based on spatial proximity followed by feature similarity. SuperGSeg [19] and InstanceGaussian [17] both leverage neural Gaussians to model instance-level features: SuperGSeg groups Gaussians into Super-Gaussians to facilitate language assignment, whereas InstanceGaussian directly assigns fused semantic features to each cluster. VoteSplat [11] and OpenSplat3D [25] mitigate the pixel-level ambiguities of the direct distillation. Then, the resulting cluster graph structures support higher-level reasoning [19, 39], which per-Gaussian features cannot easily enable. However, all these methods rely on feature distillation using per-cluster learnable language embeddings. These approaches increase computational and implementation complexity with their typical three-stage pipeline: (1) training 3DGS, (2) clustering, and (3) assigning language features to the clusters and individual Gaussians. Then, by assigning the same feature to an entire cluster, fine-grained details are lost.
@@
Open-Vocabulary Feature Aggregation. Beyond cluster-based language features distillation, recent works adopt more efficient strategies for feature aggregation. For instance, Dr.Splat [12] and Occam’s LGS [3] bypass intermediate 2D supervision and clustering by directly injecting language features into 3D Gaussians, achieving fast, accurate results in a training-free regime. While these direct feature aggregation methods deliver strong runtime efficiency and segmentation accuracy, they indiscriminately propagate 2D features to every Gaussian intersected by each camera ray, disregarding scene geometry and occlusion. As a result, features from foreground objects (e.g., a vase) are erroneously assigned to background elements (e.g., the table or floor). Moreover, existing methods share two critical limitations: (i) they assign equal supervision to all Gaussians along a ray, ignoring each Gaussian’s marginal contribution to the rendered pixel, and (ii) they overlook the view-dependent noise and inconsistency in 2D language features. We address these issues with VALA, a robust and efficient training-free framework that improves segmentation through visibility-aware gating (for contribution-aligned supervision) and robust multi-view aggregation.
@@ 3 Preliminaries
Language Features Assignment via Direct Aggregation. Recent works [12, 3] proposed to directly assign 2D language features to 3D Gaussians via weighted feature aggregation. To obtain training-free 3D language feature embeddings, Kim et al[12] pool per-pixel weights $w_{i}(I,r)$, defined as in Eq. (5), using segmentation masks $M_{j}(I,r)$:
@@ 4 Method
We aim to distill language features into 3DGS under visibility constraints, to get semantically rich and view-consistent 3D embeddings. Unlike existing approaches that indiscriminately assign identical 2D features to all Gaussians along a camera ray, leading to noisy supervision and cross-view inconsistencies, with VALA, we assign only visible features.
Our pipeline is shown in Figure 2. Built on a direct feature assignment method, VALA has two complementary components to improve the assignment of 2D vision-language features to the 3D scene. First, we introduce a visibility-aware attribution mechanism to selectively assign language features to Gaussians based on their relevance in the rendered scene (Section 4.1). Second, we propose a robust cross-view consolidation strategy aggregating per-view features while suppressing inconsistent observations, yielding coherent 3D semantic embeddings (Section 4.2).
@@ 4.1 Visibility-Aware Feature Lifting
Recent works explored lifting 2D language embeddings into 3D space via differentiable rendering pipelines [3, 12]. However, existing approaches assign the same 2D language feature to all Gaussians intersected by a given pixel ray, regardless of each Gaussian’s actual contribution to the rendered pixel. As illustrated in Figure 3, when an object $O_{2}$ is occluded by another object $O_{1}$, the 2D language embedding at that pixel primarily represents the semantics of $O_{1}$. Nevertheless, a Gaussian $g_{2}$ belonging to $O_{2}$ may still be incorrectly associated with the language feature $f_{2}^{1}$ of $O_{1}$.
@@
where $\alpha_{i}(r)$ encodes coverage (i.e., how much $g_{i}$ overlaps the pixel), $T_{i}(r)$ transmittance (i.e., how much light reaches $g_{i}$ after occlusion by nearer Gaussians), and $w_{i}(r)$ measures how strongly $g_{i}$ influences the rendered sample along $r$. We name this as the Visibility of a Gaussian from a specific view. Instead of assigning this feature to all Gaussians on the ray $r$, we use a two-stage visibility-aware gate (VAG). We aggregate the weights into a per-view visibility score
@@
Why Mass then Quantile? A fixed quantile alone tightly controls cardinality but ignores how visibility mass is distributed, and under heavy tails may discard essential contributors. Conversely, mass coverage secures a target fraction of visible content but can be liberal when scores are flat. Our two-stage rule reconciles both: Stage A guarantees coverage on the relevant (floored) set, while B imposes a quantile-derived cardinality constraint $K_{q}^{s}$ that stabilizes scale across views. Practically, if $K_{q}^{s}\!\geq\!k_{\mathrm{mass}}^{\star}$, we keep the mass-coverage set unchanged; otherwise we truncate it to the top-$K_{q}^{s}$ by $w_{i}$. The gate is thus coverage-faithful, scale-adaptive, and compute-bounded.
@@ 4.2 Robust Multi-View Aggregation
SAM+CLIP preprocessing pipelines [27] yield crisp mask boundaries and per-pixel open-vocabulary embeddings, but their semantics are often viewpoint-dependent: changes in viewpoint and occlusion induce noticeable drift across views. To enforce multi-view consistency, several 3DGS-based methods first form 3D-consistent clusters, typically supervised with contrastive signals derived from SAM masks, and then assign a language embedding to each cluster [37, 19, 17, 25]. While this decoupled clustering can improve multi-view cross-view semantic consistency, it makes the pipelines’ training multi-stage and thus prolongs the training time. More critically, because clustering is still driven by noisy, view-dependent 2D cues, it does not correct the root cause, namely, upstream semantic drift, which can bias the clusters and ultimately degrade the accuracy of the final language assignments.
To address this multi-view inconsistency at source, we adopt geometric median [20, 36, 1] to robustly aggregate multi-view features by minimizing the cosine distances in feature space; unlike aggregation by weighted mean, it dampens view-dependent outliers and semantic drift.
@@
where $w_{i}(r)$ denotes the visibility weight of Gaussian $g_{i}$ from view $s$, since $r$ represents the view $s$. The Riemannian (projected) gradient of $\ell(\mathbf{f},\mathbf{z})=1-\mathbf{f}^{\top}\mathbf{z}$ on $\mathbb{S}^{d-1}$ is $\nabla_{\mathbf{z}}\ell=-\bigl{[}\mathbf{f}-(\mathbf{f}^{\top}\mathbf{z})\,\mathbf{z}\bigr{]},$ the projection of $\mathbf{f}$ onto the tangent space at $\mathbf{z}$. Compared to the Euclidean formulation in Eq. (14), this objective directly optimizes angular consensus, circumventing the scale sensitivity of Euclidean distances in high dimensions, where norm variations dominate over angular differences, and empirically leads to more stable 3D semantics (Table 3).
@@
Constant-Memory Streaming Update. While effective, solving Eq. (15) with the classical Weiszfeld algorithm requires repeated full-batch updates over all Gaussian features, which scales linearly with the number of views and becomes computationally prohibitive. To address this, we adopt a constant-memory streaming scheme inspired by online optimization [15]. Specifically, as detailed in Algorithm 1, we maintain only the current estimate $(\mathbf{z}_{i,t},W_{i,t})$, where $W_{i,t}$ is the cumulative visibility weight, and incorporate each new observation $(\mathbf{f}_{t},w_{i}^{t})$ via
@@
$\eta_{t}$ $=\frac{w_{i}^{t}}{W_{i,t}+w_{i}^{t}}\,,\qquad W_{i,t+1}=W_{i,t}+w_{i}^{t},$ (17)
where $\mathrm{Norm}(\mathbf{x})=\mathbf{x}/\|\mathbf{x}\|_{2}$ ensures $\mathbf{z}_{i,t}\in\mathbb{S}^{d-1}$ at every step. The direction $\mathbf{f}_{t}-(\mathbf{f}_{t}^{\top}\mathbf{z}_{i,t})\mathbf{z}_{i,t}$ corresponds to the tangent component that increases cosine similarity, while the adaptive step size $\eta_{t}$ ensures each sample contributes proportionally to its visibility. Under standard stochastic approximation assumptions (bounded variance and diminishing step sizes), $\mathbf{z}_{i,t}$ converges to a stationary point of Eq. (15) at rate $\mathcal{O}(1/\sqrt{W_{i,t}})$.
@@ 5.1 Experimental setup
Datasets.
We evaluate on the two reference datasets for this task: LERF-OVS [27] and ScanNet-v2 [5]. LERF-OVS is derived from the LERF dataset of Kerr et al[14], where we evaluate open-vocabulary object selection in both 2D and 3D. For the 2D evaluation, we follow the protocol of LERF [14]. For the 3D evaluation, we follow OpenGaussian [37]. On ScanNet, we evaluate 3D semantic segmentation. All results in each table follow the same protocol. Exhaustive details can be found in the Appendix.
@@ 5.2 Analysis on LeRF-OVS dataset
Quantitatives In 3D. The advantage of our method becomes more pronounced in 3D, with ours exceeding all baselines by a notable margin. The second best, CAGS [33], is a substantial 7.2 absolute mIoU points behind. The scene-level analysis reveals that our approach leads in Ramen, Teatime, and Waldo_Kitchen, and ranks second in Figurines, behind VoteSplat [11] due to its specialized multi-view voting. The gains are especially significant in large, cluttered environments (Teatime, Waldo_Kitchen), where our contribution-aware aggregation better preserves semantics through severe occlusions.
The strong 3D consistency of our method contrasts with approaches like LangSplat and LEGaussian [30], whose high 2D accuracy does not translate to 3D performance, likely due to their lack of explicit handling of per-ray contribution and occlusion. Similarly, the post-hoc clustering methods OpenGaussian [37] and SuperGSeg [19] show moderate 3D improvements but remain sensitive to the upstream semantic drift, limiting their robustness. Our performance relative to Occam’s LGS (baseline) is noteworthy: while both adopt streaming updates, our visibility-guided feature attribution yields much higher mIoU and mAcc in 3D, highlighting the effectiveness of improving the semantic assignment at the feature aggregation stage rather than solely relying on memory-efficient training.
Qualitatives in 3D. We show visual 3D results in Figure 4. Existing approaches, such as InstanceGaussian [17], frequently fail by retrieving incorrect objects across multiple scenes. This can be attributed to their reliance on appearance–semantic joint representations, which struggle to distinguish small objects with visually similar appearances. Clustering-based methods struggle with multiple nearby instances. For example, querying for “knife”, OpenGaussian [37] and InstanceGaussian [17] detect only one out of five knives, whereas Dr.Splat [12] and Occam’s LGS [3] identify all knives but produce indistinct boundaries. In contrast, ours successfully localizes all knives with accurate and sharp delineations. Our approach also demonstrates robustness on challenging small-object queries, such as “Kamaboko” and “egg” in the Ramen scene. These targets lie within a heavily cluttered context (a bowl of ramen), making them particularly difficult to isolate. Competing methods [37, 12, 17] fail to recognize these objects, while Occam’s LGS correctly retrieves them but with blurred contours. By comparison, ours produces precise boundaries and accurately captures fine object structures. Similar improvements are observed in the “Spatula” query, further illustrating that our visibility-aware gating not only mitigates occlusion effects but also enables the recovery of fine-grained details in complex scenes.
@@
Qualitatives. Qualitative comparisons are presented in Figure 5. In the large and complex second room, our method accurately predicts the wall behind the bed (bed in orange), a structure often misclassified by others. In the smaller but more occluded third scene, our method also demonstrates superior 3D segmentation, better capturing challenging objects such as the central table. This ability to recover occluded and fine-scale geometry is particularly beneficial for downstream applications such as 3D object localization. Overall, the qualitative results support the quantitative improvements, highlighting both the robustness and effectiveness of our proposed framework.

### v2

@@ Abstract
Recently, distilling open-vocabulary language features from 2D images into 3D Gaussians has attracted significant attention. Although existing methods achieve impressive language-based interactions with 3D scenes, we observe two fundamental issues: background Gaussians, which contribute negligibly to a rendered pixel, receive the same feature as the dominant foreground ones, and multi-view inconsistencies due to view-specific noise in language embeddings. We introduce Visibility-Aware Language Aggregation (VALA), a lightweight yet effective method that computes marginal contributions for each ray and applies a visibility-aware gate to retain only visible Gaussians. Moreover, we propose a streaming weighted geometric median in cosine space to merge noisy multi-view features. Our method yields a robust, view-consistent language feature embedding in a fast and memory-efficient manner. VALA improves open-vocabulary localization and segmentation across reference datasets, consistently surpassing existing works. The source code is available on VALA.
@@ 1 Introduction
Understanding 3D scenes is essential for interacting with the environment in robotic navigation [2, 23], autonomous driving [8, 32], and augmented reality [17, 10]. Traditional approaches, however, are constrained to a fixed set of object categories defined at training time [27, 4, 35], limiting their applicability to open-world scenarios. Thanks to recent advances in vision-language models [30, 11], open-vocabulary methods [9, 42, 24] enable querying and interacting with 3D scenes through natural language, and recognizing unseen object categories without requiring retraining.
While classical 3D understanding methods operate on point clouds or meshes derived from 3D sensors, recent neural scene representations, such as NeRFs [22] and 3D Gaussian Splatting (3DGS) [14], have emerged as a compelling alternative. They not only enable high-quality rendering from novel viewpoints but also facilitate semantic reasoning, as appearance and geometry are encoded jointly. Thus, open-vocabulary reasoning has recently been grounded in neural 3D scene representations [15, 28], enabling new semantic interactions in 3D environments. Initially explored with NeRFs [15, 7], the efficiency and explicit nature of 3DGS simplified the integration of semantic features, contributing to its widespread adoption [28, 38, 13, 3].
@@
Despite numerous recent advances [13, 12, 18, 34], current approaches suffer from an inherent limitation: they assign 2D features indiscriminately to all Gaussians along a camera ray, disregarding scene geometry and occlusion relationships. Consequently, features originating from foreground objects (e.g., a vase) are incorrectly propagated to background structures (e.g., the supporting table or floor), leading to substantial degradation in open-vocabulary recognition accuracy.
Furthermore, when lifted into 3D, 2D features exhibit multi-view inconsistencies. The same object may produce divergent feature representations across different viewpoints, a phenomenon known as semantic drift [15]. Current methods address this by promoting cross-view consistency through 3D-consistent clustering and contrastive objectives derived from SAM masks [38, 20, 18, 26]. Nevertheless, such strategies generally require extensive per-scene optimization, and their heavy reliance on noisy, view-dependent 2D cues often undermines cluster reliability.
@@
  • We propose a robust aggregation strategy for the 2D features using the streaming cosine median, thereby improving multi-view consistency.
@@ 2 Related works
Other methods first group 3D Gaussians or points into semantically meaningful clusters, typically corresponding to objects or parts, and then assign a language feature to each cluster as a whole [20, 38, 29, 18, 26, 12]. These methods introduce an explicit discrete grouping step as a form of prior semantic structuring: OpenGaussian [38] performs coarse-to-fine clustering based on spatial proximity followed by feature similarity. SuperGSeg [20] and InstanceGaussian [18] both leverage neural Gaussians to model instance-level features: SuperGSeg groups Gaussians into Super-Gaussians to facilitate language assignment, whereas InstanceGaussian directly assigns fused semantic features to each cluster. VoteSplat [12] and OpenSplat3D [26] mitigate the pixel-level ambiguities of the direct distillation. Then, the resulting cluster graph structures support higher-level reasoning [20, 40], which per-Gaussian features cannot easily enable. However, all these methods rely on feature distillation using per-cluster learnable language embeddings. These approaches are computationally expensive and highly sensitive to noise or outliers in the preprocessed feature maps, since the language features are optimized directly in Euclidean space. As a result, even minor errors in the input features can propagate through the model, leading to inconsistent or inaccurate semantic representations, particularly in complex or cluttered scenes.
@@
Open-Vocabulary Feature Aggregation. Beyond cluster-based language features distillation, recent works adopt more efficient strategies for feature aggregation. For instance, Dr.Splat [13] and Occam’s LGS [3] bypass intermediate 2D supervision and clustering by directly injecting language features into 3D Gaussians, achieving fast, accurate results in a training-free regime. While these direct feature aggregation methods deliver strong runtime efficiency and segmentation accuracy, they indiscriminately propagate 2D features to every Gaussian intersected by each camera ray, disregarding scene geometry and occlusion. As a result, features from foreground objects (e.g., a vase) are erroneously assigned to background elements (e.g., the table or floor). Moreover, existing methods share two critical limitations: (i) they assign equal supervision to all Gaussians along a ray, ignoring each Gaussian’s marginal contribution to the rendered pixel, and (ii) they overlook the view-dependent noise and inconsistency in 2D language features. We address these issues with VALA, a robust and efficient training-free framework that improves open-vocabulary grounding through visibility-aware gating (for contribution-aligned supervision) and robust multi-view aggregation.
@@ 3 Preliminaries
Language Features Assignment via Direct Aggregation. Recent works [13, 3] proposed to directly assign 2D language features to 3D Gaussians via weighted feature aggregation. To obtain training-free 3D language feature embeddings, Kim et al. [13] pool per-pixel weights $w_{i}(I,r)$, defined as in Eq. (5), using segmentation masks $M_{j}(I,r)$:
@@ 4 Method
We aim to distill language features into 3DGS under visibility constraints, to get semantically rich and view-consistent 3D embeddings. Existing approaches that indiscriminately assign identical 2D features to all Gaussians along a camera ray, which leads to noisy supervision and cross-view inconsistencies. With VALA, we assign only visible features.
Our pipeline is shown in Figure 2. Built on a direct feature assignment method, VALA has two complementary components to improve the assignment of 2D vision-language features to the 3D scene. First, we introduce a visibility-aware attribution mechanism to selectively assign language features to Gaussians based on their relevance in the rendered scene (Section 4.1). Second, we propose a robust cross-view consolidation strategy to aggregate per-view features while suppressing inconsistent observations, yielding coherent 3D semantic embeddings (Section 4.2).
@@ 4.1 Visibility-Aware Feature Lifting
Recent works explored lifting 2D language embeddings into 3D space via differentiable rendering pipelines [3, 13]. However, existing approaches assign the same 2D language feature to all Gaussians intersected by a given pixel ray, regardless of each Gaussian’s actual contribution to the rendered pixel. As illustrated in Figure 3, when an object $O_{2}$ is occluded by another object $O_{1}$, the 2D language embedding at that pixel primarily represents the semantics of $O_{1}$. Nevertheless, a Gaussian $g_{2}$ belonging to $O_{2}$ may still be incorrectly associated with the language feature of $O_{1}$.
@@
where $\alpha_{i}(r)$ encodes coverage (i.e., how much $g_{i}$ overlaps the pixel), $T_{i}(r)$ represents transmittance (i.e., how much light reaches $g_{i}$ after occlusion by nearer Gaussians), and $w_{i}(r)$ measures how strongly $g_{i}$ influences the rendered sample along $r$. We name this as the Visibility of a Gaussian from a specific view. Instead of assigning this feature to all Gaussians on the ray $r$, we use a two-stage visibility-aware gate (VAG). We aggregate the weights into a per-view visibility score
@@
Why Mass then Quantile? A fixed quantile alone tightly controls cardinality but ignores how visibility mass is distributed, and under heavy tails may discard essential contributors. Conversely, mass coverage secures a target fraction of visible content but can be liberal when scores are flat. Our two-stage rule reconciles both: Stage A guarantees coverage on the relevant (floored) set, while B imposes a quantile-derived cardinality constraint $K_{q}^{s}$ that stabilizes scale across views. Practically, if $K_{q}^{s}\!\geq\!k_{\mathrm{mass}}^{\star}$, we keep the mass-coverage set unchanged; otherwise we truncate it to the top-$K_{q}^{s}$ by $w_{i}$. The gate is thus coverage-faithful and scale-adaptive.
@@ 4.2 Robust Multi-View Aggregation
SAM+CLIP preprocessing pipelines [28] yield crisp mask boundaries and per-pixel open-vocabulary embeddings, but their semantics are often viewpoint-dependent: changes in viewpoint and occlusion induce noticeable drift across views. To enforce multi-view consistency, several 3DGS-based methods first form 3D-consistent clusters, typically supervised with contrastive signals derived from SAM masks, and then assign a language embedding to each cluster [38, 20, 18, 26]. While this decoupled clustering can improve multi-view semantic consistency, it makes the pipelines’ training multi-stage, thus prolonging the training time. More critically, because clustering is still driven by noisy, view-dependent 2D cues, it does not correct the root cause, namely, upstream semantic drift, which can bias the clusters and ultimately degrade the accuracy of the final language assignments.
To address this multi-view inconsistency at source, we adopt geometric median [21, 37, 1] to robustly aggregate multi-view features by minimizing the cosine distances in feature space. Unlike aggregation by weighted mean, it dampens view-dependent outliers and semantic drift.
@@
where $w_{i}(r)$ denotes the visibility weight of Gaussian $g_{i}$ from view $s$, since $r$ represents the view $s$. The gradient of $\ell(\mathbf{f},\mathbf{z})=1-\mathbf{f}^{\top}\mathbf{z}$ on $\mathbb{S}^{d-1}$ is $\nabla_{\mathbf{z}}\ell=-\bigl[\mathbf{f}-(\mathbf{f}^{\top}\mathbf{z})\,\mathbf{z}\bigr],$ the projection of $\mathbf{f}$ onto the tangent space at $\mathbf{z}$. Compared to the Euclidean formulation in Eq. (14), this objective directly optimizes angular similarity, circumventing the scale sensitivity of Euclidean distances in high dimensions, where norm variations dominate over angular differences, and empirically leads to more stable 3D semantics (Table 3).
@@
Constant-Memory Streaming Update. While effective, solving Eq. (15) with the classical Weiszfeld algorithm [6] requires repeated full-batch updates over all Gaussian features, which scales linearly with the number of views and becomes computationally prohibitive in practice. To address this, we adopt a constant-memory streaming scheme inspired by online optimization [16]. Specifically, as detailed in Algorithm 1, we maintain only the current estimate $(\mathbf{z}_{i,t},W_{i,t})$, where $W_{i,t}$ is the cumulative visibility weight, and incorporate each new observation $(\mathbf{f}_{t},w_{i}^{t})$ via
@@
$\eta_{t}$ $=\frac{w_{i}^{t}}{W_{i,t}+w_{i}^{t}}\,,\qquad W_{i,t+1}=W_{i,t}+w_{i}^{t},\vskip-5.69054pt$ (17)
where $\mathrm{Norm}(\mathbf{x})=\mathbf{x}/\|\mathbf{x}\|_{2}$ projects $\mathbf{z}_{i,t}$ onto the unit sphere $\mathbb{S}^{d-1}$. The update direction $\mathbf{f}_{t}-(\mathbf{f}_{t}^{\top}\mathbf{z}_{i,t})\mathbf{z}_{i,t}$ lies in the tangent space and increases cosine similarity, while the adaptive step size $\eta_{t}$ weights each sample according to its visibility. Under standard stochastic approximation assumptions, $\mathbf{z}_{i,t}$ converges to a stationary point of Eq. (15) at rate $\mathcal{O}(1/\sqrt{W_{i,t}})$.
@@ 5.1 Experimental setup
Datasets. We evaluate on the two reference datasets for this task: LERF-OVS [28] and ScanNet-v2 [5]. LERF-OVS is derived from the LERF dataset of Kerr et al. [15], where we evaluate open-vocabulary object selection in both 2D and 3D. For the 2D evaluation, we follow the protocol of LERF [15]. For the 3D evaluation, we follow OpenGaussian [38]. On ScanNet, we evaluate 3D semantic segmentation. Previous evaluation protocols [38, 18] freeze the growth of 3D Gaussians, which degrades photometric fidelity. In contrast, we allow full optimization of the 3D Gaussians, resulting in misalignment between the optimized Gaussians and the ground-truth point cloud. We therefore adapt the evaluation protocol in [13] by propagating pseudo ground-truth labels to the Gaussians. Details are provided in the Appendix B.
@@ 5.2 Analysis on LeRF-OVS dataset
Quantitatives In 3D. The advantage of our method becomes more pronounced in 3D, with ours exceeding all baselines by a notable margin. The second best, CAGS [34], is a substantial 7.2 absolute mIoU points behind. The scene-level analysis reveals that our approach leads in Ramen, Teatime, and Waldo_Kitchen, and ranks second in Figurines, behind VoteSplat [12] due to its specialized multi-view voting. The gains are especially significant in large, cluttered environments (Teatime, Waldo_Kitchen), where our contribution-aware aggregation better preserves semantics despite severe occlusions.
The strong 3D consistency of our method contrasts with approaches like LangSplat and LEGaussian [31], whose high 2D accuracy does not translate to 3D performance, likely due to their lack of explicit handling of per-ray contribution and occlusion. Similarly, the post-hoc clustering methods OpenGaussian [38] and SuperGSeg [20] exhibit moderate 3D improvements but remain sensitive to upstream semantic drift, thereby limiting their robustness. Our performance relative to Occam’s LGS (baseline) is noteworthy: while both adopt streaming updates, our visibility-guided feature attribution yields much better performance in 3D, highlighting the effectiveness of improving the semantic assignment at the feature aggregation stage rather than solely relying on memory-efficient training.
Qualitatives in 3D. We show visual 3D results in Figure 4. Existing approaches, such as InstanceGaussian [18], frequently fail by retrieving incorrect objects across multiple scenes. This can be attributed to their reliance on appearance–semantic joint representations, which struggle to distinguish small objects with visually similar appearances. Clustering-based methods struggle with multiple instances that are closely related. For example, querying for “knife”, OpenGaussian [38] and InstanceGaussian [18] detect only one out of five knives, whereas Dr.Splat [13] and Occam’s LGS [3] identify all knives but produce indistinct boundaries. In contrast, ours successfully localizes all knives with accurate and sharp delineations. Our approach also demonstrates robustness on challenging small-object queries, such as “Kamaboko” and “egg” in the Ramen scene. These targets lie within a heavily cluttered context (a bowl of ramen), making them particularly difficult to isolate. Competing methods [38, 13, 18] fail to recognize these objects, while Occam’s LGS correctly retrieves them but with blurred contours. By comparison, ours produces precise boundaries and accurately captures fine object structures. Similar improvements are observed in the “Spatula” query, further illustrating that our visibility-aware gating not only mitigates occlusion effects but also enables the recovery of fine-grained details in complex scenes.
@@
Qualitatives. Qualitative comparisons are presented in Figure 5. In the large and complex second room, our method accurately predicts the wall behind the bed (bed in orange), a structure often misclassified by others. In the smaller but more occluded third scene, our method also demonstrates superior 3D segmentation, capturing challenging objects such as the central table more effectively. This ability to recover occluded and fine-scale geometry is particularly beneficial for downstream applications such as 3D object localization. Overall, the qualitative results support the quantitative improvements, highlighting both the robustness and effectiveness of our proposed framework.